# Color it, Code it, Cancel it: $k$-local dynamical decoupling from classical additive codes

### v1

@@ IV.4 Selective DD sequences
We now apply our formalism to selectively cancel only particular interactions, rather than to completely decouple the system from the environment or suppress all residual interactions. To do so, we assume that the system Hamiltonian $H_{S}$ can be decomposed into two parts:
@@
where $H_{S}^{\parallel}$ contains the interaction terms we wish to preserve and $H_{S}^{\perp}$ contains the terms we aim to suppress. This task is particularly relevant for applications such as Hamiltonian simulation, where it is desirable to selectively engineer specific interactions [29].
Based on the connection between DD sequences and error-detection codes, we can construct a simple algorithm that finds the DD group performing this task. Our method explicitly searches for generators of the DD group that suppresses the undesired terms in $H_{S}^{\perp}$ while preserving the desired terms in $H_{S}^{\parallel}$.
As a first step, we identify a set of generators that commute with all terms in $H_{S}^{\parallel}$. This is efficiently achieved by representing each Pauli string in $H_{S}^{\parallel}$ as a binary vector in $\mathbb{F}_{2}^{2n}$ via the mapping
$\vec{P}=\left(\prod_{i=1}^{n}X_{i}^{x_{i}}Z_{i}^{z_{i}}\right)~\to~\begin{pmatrix}\vec{x}~|~\vec{z}\end{pmatrix}^{T}\quad\text{ with }\vec{x},\vec{z}\in\mathbb{F}_{2}^{n}.$ (34)
In this representation, the commutation relation between two Pauli strings $\bf{u}$ and $\bf{v}$ is determined by their symplectic inner product
$\bf{u}^{T}\Omega\bf{v},\quad\text{ where }\quad\Omega=\begin{pmatrix}0_{n\times n}&I_{n\times n}\\ I_{n\times n}&0_{n\times n}\end{pmatrix}$. (35)
@@
Given the bipartite graph, we numerically identify the minimal set of generators $\{\bf{n}_{i}\}$ that preserves $H_{S}^{\parallel}$ while decoupling $H_{S}^{\perp}$. This reduces to the classical minimum set cover problem: finding the smallest subset of $\{\bf{n}_{i}\}$ such that every $\bf{h}_{i}^{\perp}$ anticommutes with at least one selected generator. While set cover is generally a NP-hard problem, this problem remains tractable in simple cases, where we assume the chromatic number $\chi_{I}$ is small, an assumption typically valid for planar devices with short range interactions.
We emphasize that this algorithm is primarily a proof of concept rather than a fully optimized solution. Nonetheless, it shows that the error-detection perspective can drastically reduce the search space by focusing only on the generators of the decoupling group instead of the entire sequence, generally enabling a more efficient and automated construction of tailored decoupling sequences. This viewpoint could also potentially benefit machine learning–based approaches to decoupling design, as explored in recent work [33, 34].
@@ V Applications to quantum computing
In the previous section, we linked the problem of designing time-efficient DD sequences for arbitrary $k$-local Hamiltonians to classical error detection codes. Here, we explicitly construct DD sequences for 2-local and 3-local Hamiltonians, tailored to superconducting and spin qubits, using known constructions of classical additive codes. As an example of Hamiltonian engineering, we also introduce a sequence that enables the simulation of the Kitaev’s honeycomb model in spin qubit platforms.
Figure 5: Colored graphs of common superconducting qubit architectures. We assume a noise model where each qubit experiences both dephasing ($Z$ errors) and relaxation ($X$ and $Y$ errors), along with residual $ZZ$ interactions between neighbors and $ZZZ$ interactions among any three consecutive qubits. The colorings also apply when residual $ZZ$ interactions extend to next-nearest neighbors. (a) Heavy-hex lattice with $\chi_{I}=6$ colors. (b) Square lattice with $\chi_{I}=6$ colors. (c) The resulting quotient graph is identical for both architectures.
@@ V.1.1 Suppressing $ZZZ$ coupling
In superconducting platforms, suppressing residual $ZZ$ and $ZZZ$ couplings between neighboring qubits is a key challenge for scalability, as these interactions are a major source of coherent errors [61, 62, 63].
In this section, we show that by combining the graph coloring procedure introduced earlier with binary Reed–Muller codes, one can construct efficient DD sequences. Specifically, we introduce a bang-bang control sequence with circuit depth scaling as $O(\chi_{I})$. In Appendix B, we extend this result to bounded-strength control, yielding a DD sequence with depth scaling as $O(\chi_{I}\log\chi_{I})$.
As a concrete example, Fig. 5(a) and Fig. 5(b) show the color partitions using $\chi_{I}=6$ for the heavy-hex lattice and the square lattice, used by IBM [64] and Google [65] respectively. In this section, we assume a noise model in which each qubit is subject to both dephasing and relaxation, along with residual $ZZ$ interactions between neighboring qubits and $ZZZ$ interactions across any three consecutive qubits. The presented color partitions also apply when considering residual $ZZ$ interactions between both nearest and next-nearest neighbors. Fig. 5(c) shows the corresponding quotient graph induced by the color partition. We emphasize that both the heavy-hex and square lattices yield the same quotient graph under this partitioning. In the following, we briefly review the Reed–Muller code and outline how to construct tailored DD sequences for superconducting qubits based on it.
The Reed-Muller code is one of the oldest and simplest families of linear binary codes with well-understood parameters and dual structure [66]. A Reed-Muller code $\rm RM(r,m)$ is a binary code with parameters
@@
$\gamma_{i}\to f_{\gamma_{i}}(\bm{x})=x_{i_{1}}x_{i_{2}}\dots x_{i_{s}},$ (39)
where $\bm{x}=(x_{1},\dots,x_{m})\in\mathbb{F}_{2}^{m}$, $0\leq s\leq r$, and the total degree of $f_{g_{i}}$ is at most $r$. There are exactly
@@
of such linearly independent monomials, corresponding to the dimension of the code. Each generator $\gamma_{i}$ is a binary vector of length $2{{}^{m}}$ obtained by evaluating the associated polynomial $f_{\gamma_{i}}(\bm{x})$ at all $2^{m}$ possible input $\bm{x}\in\mathbb{F}_{2}^{m}$.
@@
If we replace each codeword in $\mathcal{C}$ by substituting 0 with the identity $I$ and $1$ with $X$, then Theorem IV.1 ensures that the resulting DD sequence removes all $Z$-type interactions with weight up to $d^{\perp}-1=3$.
The generators of $\rm RM(1,m)$ correspond to all linear Boolean functions over $\mathbb{F}_{2}^{m}$, i.e., the constant function $f_{\gamma_{0}}=1$ and the monomials $f_{\gamma_{i}}=x_{i}$ with $i=1,\dots,m$. As an example, Table 1(a) lists the generators of $\rm RM(1,3)$ as rows, corresponding to the chromatic number $\chi_{I}=2^{3}=8$. The generator $\gamma_{0}$ (first row) associated with the constant function $f_{\gamma_{0}}=1$ can detect all single-qubit $Z_{i}$ terms as well as any three-qubit $Z_{i}Z_{j}Z_{k}$ terms. In addition, the other generators $\gamma_{i\geq 1}$ detect two-qubit $Z_{i}Z_{j}$ interactions between qubits with distinct colors. Importantly, since the heavy-hex and square lattices in Fig. 5 require only $\chi_{I}=6$ colors, the DD sequences derived from Table 1(a) is directly applicable to these architectures.
@@ X X I I X X I
The DD sequence constructed from the $\rm RM(1,m)$ code using the mapping $0\to I$ and $1\to X$ can only suppresses $Z$-type interactions. However, to universally suppress all single-qubit terms in addition to residual $Z_{i}Z_{j}$ and $Z_{i}Z_{j}Z_{k}$ couplings, a simple modification suffices. The key idea is that the mapping can be generalized by assign 0 to $I$ or $Z$, and $1$ to either $X$ or $Y$. The result decoupling sequence still removes all $Z$-type interactions of weight up to three.
To universally decouple single-qubit terms, we require that for each color class the support of the generators must include at least two distinct Pauli operators from $\{X,Y,Z\}$. In other words, each column in Table 1(a) must contain at least two different Pauli matrices. This condition can always be satisfied for the generators of $\rm RM(1,m)$ by appropriately applying the substitution rule discussed above. Table 1(b) provides an example of this modified substitution applied to Table 1(a). Consequently, the length $L$ of the decoupling sequence constructed based on the Reed-Muller codes is given by
@@
In practical terms, this means that for the near-term architectures shown in Fig. 5, a tailored bang-bang decoupling sequence of length $L=16$ is sufficient to protect superconducting qubit systems against interactions up to third order. In contrast, as we will show later, using a universal (non-tailored) DD sequence that suppresses all third-order interactions requires a significantly longer sequence of length $L=64$, a fourfold increase.
@@ V.1.2 Tailored sequence to only $ZZ$ couplings
Our DD sequence based on Reed-Muller codes can universally suppress single-qubit terms, as well as two-qubit $Z_{i}Z_{j}$ and three-qubit $Z_{i}Z_{j}Z_{k}$ interactions. However, if three-qubit interactions are negligible or absent, we can improve further the time-efficiency of the sequence. This improvement originates from two factors. First, restricting to two-body interactions reduces the chromatic number $\chi_{I}$ of the interaction graph. Second, we can drop one generator from the $\rm RM(1,m)$ code, effectively halving the length of the decoupling sequence.
The generator we remove is $\gamma_{0}$ which, as previously discussed, detects all weight-three $Z$-type terms. As a consequence, the final color $C_{2^{m}}$ (corresponding to the last column of Table 1(b)) can no longer detect single-qubit $Z$ terms. Therefore, we must also remove this color. This procedure is equivalent to constructing a punctured Reed-Muller code. However, since we also discard $\gamma_{0}$, the resulting code is a modified punctured $\rm RM(1,m)$ code with dual distance $d^{\perp}=3$ instead of $d^{\perp}=4$. Consequently, the length of the DD sequence is
@@
The resulting decoupling sequence universally suppresses single-qubit terms and two-qubit $Z_{i}Z_{j}$ interactions, while requiring one fewer generator at the cost of one fewer color. In Table 1(c), we present the generators of such decoupling sequence generated from Table 1(b) by removing the first row and the last column. This modified punctured $\rm RM(1,m)$ construction is, up to minor adjustments, equivalent to the single-axis Hadamard construction introduced in Ref. [27].
This means that to protect superconducting qubit systems with the architecture shown in Fig. 5 against both nearest-neighbor and next-nearest-neighbor $ZZ$ couplings, a tailored bang-bang DD sequence of length $L=8$ suffices. In contrast, as we will show later, a universal (non-tailored) DD sequence that suppresses all interactions up to the same order requires a significantly longer sequence of length $L=32$, again a fourfold increase.
In this section, we have shown how to construct tailored bang-bang DD sequences for removing residual interactions in superconducting qubits by using binary Reed–Muller codes. In Appendix B, we extend this construction to bounded-strength control sequences. In the next section, we turn to the spin-qubit platform and introduce tailored decoupling sequences based on quaternary codes.
@@ V.2 Exchange interactions in spin qubits
Figure 6: Colored trilinear array architecture in the spin-qubit platform. We assume a noise model where each qubit undergoes both dephasing and relaxation, with Heisenberg exchange between neighbors and scalar chirality interactions among any three qubits forming a triangle. (a) The color partition $C[I_{\rm Dev}]$ uses $\chi_{I}=4$ colors, tailored to this interaction model. (b) The resulting quotient graph is a square with diagonal links.
@@ II II II $Z$ $Y$ $X$ $X$ $Z$ $Y$ $Z$ $Y$ $X$ $Z$ $Y$ $X$
Table 2: Universal suppression of two-local and Heisenberg exchange interactions. These tables illustrate how to derive DD groups from additive projective geometry codes based on a 1-spread in $\mathrm{PG}(3,2)$. (a) Left: A (complete) 1-spread in $\mathrm{PG}(3,2)$, where each column represents a line described by two projective points in $\mathbb{F}_{2}^{4}$. Center: The corresponding projective points in $\mathrm{PG}(3,4)$ constructed from the 1-spread in $\rm PG(3,2)$ using Eq. 53. Right: The generators of a DD group that suppresses universally suppress two-local interactions constructed from projective points in $\rm PG(3,4)$. (b) Generators of DD group tailored for spin-qubit systems based on a 1-spread in $\mathrm{PG}(3,2)$. Each projective point in $\mathrm{PG}(3,4)$ is expanded into its three equivalent representatives in $\mathbb{F}_{4}^{4}$, tripling the number of color classes.
Dense spin-qubit architectures are challenged by residual exchange interaction [57, 67, 56] that cannot be completely turned off. The dominant residual interactions for system with small spin-orbit interaction are the two-qubit Heisenberg exchange interactions
@@
and the scalar chirality interaction [68] (for perpendicular magnetic fields)
@@
The class of three-local Hamiltonians is particularly important, as it enables the implementation of single-step, high-fidelity three-qubit gates [56, 69, 70]. As a concrete example, we consider a trilinear array architecture [71] designed to support such gates, as illustrated in Fig. 6(a) [56]. We assume a noise model in which each qubit experiences both dephasing and relaxation, along with Heisenberg exchange interactions between neighboring qubits and scalar chirality interactions among any trio forming a triangle. Under this model, the array admits a coloring with $\chi_{I}=4$ colors. The resulting quotient graph is shown in Fig. 6(b).
To construct a tailored DD sequence eliminating all interactions up to weight three, we can generate an orthogonal array by using the projective geometry code $\mathcal{C}$ and its dual (also known as the simplex code) as $\mathcal{C}^{\perp}$. The simplex code is chosen because it enables the large number of colors $\chi_{I}$ with minimal number of generators, resulting in a highly time-efficient sequence.

### v2

@@ IV.4 Selective DD sequences
We now apply our formalism to selectively suppress only particular interactions, rather than to completely decouple the system from the environment or freeze the system dynamic. To do so, we assume that the system Hamiltonian $H_{S}$ can be decomposed into two parts:
@@
where $H_{S}^{\parallel}$ contains the interaction terms we wish to preserve and $H_{S}^{\perp}$ contains the terms we aim to suppress. This task is particularly relevant for applications such as Hamiltonian simulation, where it is desirable to selectively engineer specific interactions [Choi2020].
Based on the connection between DD sequences and error-detection codes, we can construct a simple algorithm that finds the decoupling group performing this task. Our method explicitly searches for generators of the DD group that suppresses the undesired terms in $H_{S}^{\perp}$ while preserving the desired terms in $H_{S}^{\parallel}$.
As a first step, we identify a set of generators that commute with all terms in $H_{S}^{\parallel}$. This is efficiently achieved by representing each Pauli string in $H_{S}^{\parallel}$ as a binary vector in $\mathbb{F}_{2}^{2n}$ via the symplectic representation
$P=\left(\prod_{i=1}^{n}X_{i}^{x_{i}}Z_{i}^{z_{i}}\right)~\to~\begin{pmatrix}\vec{x}~|~\vec{z}\end{pmatrix}^{T}\quad\text{ with }\vec{x},\vec{z}\in\mathbb{F}_{2}^{n}.$ (30)
In the symplectic representation, the commutation relation between two Pauli strings $\bf{u}$ and $\bf{v}$ is determined by their symplectic inner product
$\bf{u}^{T}\Omega\bf{v},\qquad\text{ where }\qquad\Omega=\begin{pmatrix}0_{n\times n}&I_{n\times n}\\ I_{n\times n}&0_{n\times n}\end{pmatrix}$. (31)
@@
Given the bipartite graph, we numerically identify the minimal set of generators $\{\bf{n}_{i}\}$ that preserves $H_{S}^{\parallel}$ while decoupling $H_{S}^{\perp}$. This reduces to the classical minimum set cover problem: finding the smallest subset of $\{\bf{n}_{i}\}$ such that every $\bf{h}_{i}^{\perp}$ anticommutes with at least one selected generator. While the minimal set cover is generally a NP-hard problem, this problem remains tractable in simple cases, where we assume the chromatic number $\chi_{I}$ is small, an assumption typically valid for planar devices with short range interactions. We remark that a minimum covering set does not always exist (i.e., some terms in $H^{\perp}_{S}$ may be product of terms in $H^{||}_{S}$). In such cases, one can instead employ the integer programming method introduced in Ref. [Choi2020].
We emphasize that this algorithm is primarily a proof of concept rather than a fully optimized solution. Nonetheless, it shows that the error-detection perspective can drastically reduce the search space by focusing only on the generators of the decoupling group instead of the entire sequence, generally enabling a more efficient and automated construction of tailored decoupling sequences. This viewpoint could also potentially benefit machine learning–based approaches [tong2024empirical, rahman2024learning] to designing robust DD sequences.
@@ V Applications to quantum computing
In the previous sections, we explained how to construct time-efficient DD sequences for arbitrary $k$-local Hamiltonians using tools from classical coding theory. In this section, we present concrete applications of our framework based on well-known families of classical codes. We postpone the discussion of sequence robustness against pulse imperfections to Sec. VI. In Sec. V.1, we show how to construct DD sequences tailored to superconducting qubit architectures using Reed–Muller codes. In Sec. V.2, we develop DD sequences that suppress general two- and three-local errors using projective geometry codes, and demonstrate how to adapt these constructions to spin-qubit platforms. Finally, in Sec. V.3, we show how our framework can be used to simulate Kitaev’s honeycomb model starting from the isotropic exchange Hamiltonian native to spin-qubit platforms.
Figure 5: Colored graphs of common superconducting qubit architectures. We assume a noise model where each qubit experiences both longitudinal coherence errors ($Z$ errors) and transversal coherence errors ($X$ and $Y$ errors), along with residual $ZZ$ interactions between nearest- and next-nearest-neighbors, and $ZZZ$ interactions among any three consecutive qubits. The colorings also apply when residual $ZZ$ interactions extend to next-nearest neighbors. (a) Heavy-hex lattice with $\chi_{I}=6$ colors. (b) Square lattice with $\chi_{I}=6$ colors. (c) The resulting quotient hypergraph is identical for both architectures.
@@ V.1.1 Suppressing $ZZZ$ coupling
In superconducting platforms, suppressing residual $ZZ$ and $ZZZ$ couplings between neighboring qubits is a key challenge for scalability, as these interactions are a major source of coherent errors [Pederson2019, Menke2022, berke2022transmon]. In this section, we show that by combining the hypergraph coloring procedure introduced earlier with binary Reed–Muller codes, one can construct efficient DD sequences tailored to superconducting qubits. Specifically, we introduce a bang-bang control sequence with length scaling as $O(\chi_{I})$. In Appendix B, we extend this result to bounded-strength control, yielding a DD sequence with length scaling as $O(\chi_{I}\log\chi_{I})$.
As a concrete example, Fig. 5(a) and Fig. 5(b) show the color partitions using $\chi_{I}=6$ for the heavy-hex lattice and the square lattice, used by IBM [ibm_heavy_hex_2021] and Google [Acharya2025] respectively. In this section, we assume a noise model in which each qubit is subject to both longitudinal and transversal coherence errors, along with residual $ZZ$ interactions between neighboring qubits and $ZZZ$ interactions across any three consecutive qubits. The presented color partitions also apply when considering residual $ZZ$ interactions between both nearest and next-nearest neighbors. Fig. 5(c) shows the corresponding quotient hypergraph induced by the color partition. We emphasize that both the heavy-hex and square lattices yield the same quotient hypergraph under this partitioning. In the following, we briefly review the Reed–Muller code and outline how to construct tailored DD sequences for superconducting qubits based on it.
The Reed-Muller code is one of the oldest and simplest families of linear binary codes with well-understood parameters and dual structure [abbe2020reed]. A Reed-Muller code $\rm RM(r,m)$ is a binary code with parameters
@@
$\gamma_{i}\to f_{\gamma_{i}}(\boldsymbol{x})=x_{i_{1}}x_{i_{2}}\dots x_{i_{s}},$ (35)
where $\boldsymbol{x}=(x_{1},\dots,x_{m})\in\mathbb{F}_{2}^{m}$, $0\leq s\leq r$, and the total degree of $f_{g_{i}}$ is at most $r$. There are exactly
@@
linearly independent monomials, corresponding to the dimension of the code. Each generator $\gamma_{i}$ is a binary vector of length $2{{}^{m}}$ obtained by evaluating the associated polynomial $f_{\gamma_{i}}(\boldsymbol{x})$ at all $2^{m}$ possible bit strings $\boldsymbol{x}\in\mathbb{F}_{2}^{m}$.
@@
If we replace each codeword in $\mathcal{C}$ by substituting 0 with the identity $I$ and $1$ with $X$, then Theorem IV.1 ensures that the resulting DD sequence suppresses all $Z$-type interactions with weight up to $d^{\perp}-1=3$.
The generators of $\rm RM(1,m)$ correspond to all linear Boolean functions over $\mathbb{F}_{2}^{m}$, i.e., the constant function $f_{\gamma_{0}}=1$ and the monomials $f_{\gamma_{i}}=x_{i}$ with $i=1,\dots,m$. As an example, Table 1(a) lists the generators of $\rm RM(1,3)$ as rows, corresponding to the chromatic number $\chi_{I}=2^{3}=8$. The generator $\gamma_{0}$ (first row) associated with the constant function $f_{\gamma_{0}}=1$ can suppress all single-qubit $Z_{i}$ terms as well as any three-qubit $Z_{i}Z_{j}Z_{k}$ terms. In addition, the other generators $\gamma_{i\geq 1}$ suppress two-qubit $Z_{i}Z_{j}$ interactions between qubits with distinct colors. Importantly, since the heavy-hex and square lattices in Fig. 5 require only $\chi_{I}=6$ colors, the DD sequence derived from Table 1(a) is directly applicable to these architectures.
@@ X X I I X X I
The DD sequence constructed from the $\rm RM(1,m)$ code using the mapping $0\to I$ and $1\to X$ can only suppress $Z$-type interactions. However, to universally suppress all single-qubit terms in addition to residual $Z_{i}Z_{j}$ and $Z_{i}Z_{j}Z_{k}$ couplings, a simple modification suffices. The key idea is that the mapping can be generalized by assigning "0" to $I$ or $Z$, and "11" to either $X$ or $Y$. The resulting decoupling sequence still removes all $Z$-type interactions of weight up to three.
To universally decouple single-qubit terms, we require that the support of the generators on each color class must include at least two distinct Pauli operators from $\{X,Y,Z\}$. In other words, each column in Table 1(a) must contain at least two different Pauli matrices. This condition can always be satisfied for the generators of $\rm RM(1,m)$ by appropriately applying the substitution rule discussed above. Table 1(b) provides an example of the modified substitution applied to Table 1(a). Consequently, the length $L$ of the decoupling sequence constructed based on the Reed-Muller codes is given by
@@
In practical terms, this means that for the near-term architectures shown in Fig. 5, a tailored bang-bang decoupling sequence of length $L=16$ is sufficient to protect superconducting qubit systems against 3-local interactions. In contrast, as we will show later, using a universal DD sequence that suppresses all 3-local interactions requires a significantly longer sequence of length $L=64$, a fourfold increase.
@@ V.1.2 Tailored sequence to only $ZZ$ couplings
Our DD sequence based on Reed-Muller codes can universally suppress single-qubit terms, as well as two-qubit $Z_{i}Z_{j}$ and three-qubit $Z_{i}Z_{j}Z_{k}$ interactions. However, if three-qubit interactions are negligible or absent, we can further improve the length of the sequence. The improvement originates from two factors. First, restricting to two-body interactions reduces the chromatic number $\chi_{I}$ of the interaction hypergraph. Second, we can drop one generator from the $\rm RM(1,m)$ code, effectively halving the length of the decoupling sequence.
The generator we remove is $\gamma_{0}$ which, as previously discussed, suppresses all weight-three $Z$-type terms. As a consequence, the final color class $C_{2^{m}}$ (corresponding to the last column of Table 1(b)) can no longer suppress single-qubit $Z$ terms. Therefore, we must also remove this color class. Removing color class $C_{2^{m}}$ is equivalent to constructing a punctured Reed-Muller code. However, since we also discard $\gamma_{0}$ generator, the resulting code is a modified punctured $\rm RM(1,m)$ code with dual distance $d^{\perp}=3$ instead of $d^{\perp}=4$. Consequently, the length of the DD sequence is
@@
The resulting decoupling sequence universally suppresses single-qubit terms and two-qubit $Z_{i}Z_{j}$ interactions, while requiring one fewer generator at the cost of one fewer color. In Table 1(c), we present the generators of the decoupling sequence generated from Table 1(b) by removing the first row and the last column. This modified punctured $\rm RM(1,m)$ construction is, up to minor adjustments, equivalent to the single-axis Hadamard construction introduced in Ref. [brown2024efficient].
This means that to protect superconducting qubit systems with the architecture shown in Fig. 5 against both nearest-neighbor and next-nearest-neighbor $ZZ$ couplings, a tailored bang-bang DD sequence of length $L=8$ suffices. In contrast, as we will show later, a universal DD sequence that suppresses all interactions up to the same range requires a significantly longer sequence of length $L=32$, again a fourfold increase.
In this section, we have shown how to construct tailored bang-bang DD sequences for suppressing residual interactions in superconducting qubits by using binary Reed–Muller codes. In Appendix B, we extend this construction to bounded-strength control sequences. In the next section, we turn to the spin-qubit platform and introduce tailored decoupling sequences based on quaternary codes.
@@ V.2 Exchange interactions in spin qubits
Figure 6: Colored trilinear array architecture in the spin-qubit platform. We assume a noise model where each qubit undergoes both longitudinal and transversal coherent error, with Heisenberg exchange between neighbors and scalar chirality interactions among any three qubits forming a triangle. (a) The color partition $C[I_{\rm Dev}]$ uses $\chi_{I}=4$ colors, tailored to this interaction model. (b) The resulting quotient hypergraph is a square with diagonal links.
@@ II II II $Z$ $Y$ $X$ $X$ $Z$ $Y$ $Z$ $Y$ $X$ $Z$ $Y$ $X$
Table 2: Universal suppression of two-local and Heisenberg exchange interactions. These tables illustrate how to derive DD groups from additive projective geometry codes based on a 1-spread in $\mathrm{PG}(3,2)$. (a) Left: A (complete) 1-spread in $\mathrm{PG}(3,2)$, where each column represents a line described by two projective points in $\mathbb{F}_{2}^{4}$. Center: The corresponding projective points in $\mathrm{PG}(3,4)$ constructed from the 1-spread in $\rm PG(3,2)$ using Eq. (49). Right: The additive generators of a DD group that suppresses universally suppress two-local interactions constructed from projective points in $\rm PG(3,4)$. (b) Generators of DD group tailored for spin-qubit systems based on a 1-spread in $\mathrm{PG}(3,2)$. Each projective point in $\mathrm{PG}(3,4)$ is expanded into its three equivalent representatives in $\mathbb{F}_{4}^{4}$, tripling the number of color classes.
Dense spin-qubit architectures are challenged by residual exchange interaction [Geyer2024, bosco2024exchangeonlyspinorbitqubitssilicon, nguyen2025single] that cannot be completely turned off. The dominant residual interactions for spin-qubit systems with small spin-orbit interaction are the two-qubit Heisenberg exchange interactions
@@
and the scalar chirality interaction [Sen1995] (for perpendicular magnetic fields)
@@
The class of three-local Hamiltonians is particularly important, as it enables the implementation of single-step, high-fidelity three-qubit gates [nguyen2025single, Gullans2019, Jiaan2024]. As a concrete example, we consider a trilinear array architecture [john2024two] designed to support such gates, as illustrated in Fig. 6(a) [nguyen2025single]. We assume a noise model in which each qubit experiences both longitudinal and transversal coherent error, along with Heisenberg exchange interactions between neighboring qubits and scalar chirality interactions among any trio forming a triangle. Under this interaction model, the array admits a color partition with $\chi_{I}=4$ colors. The resulting quotient hypergraph is shown in Fig. 6(b).
To construct a tailored DD sequence suppressing all interactions up to weight three, we can generate an orthogonal array using the projective geometry code $\mathcal{C}$ and its dual (also known as the simplex code) as $\mathcal{C}^{\perp}$. The code is chosen because it enables large chromatic number $\chi_{I}$ using minimal number of generators, resulting in a highly time-efficient sequence.